OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y = 4/9x – 2 that passes through the point (4,3)?

OpenStudy (anonymous):

In Statistical Mathematics the product of the gradients of two lines is -1

OpenStudy (anonymous):

Use this : Y2-Y=m(X2-X) In there : X2=4 Y2=3 Got it ?:)

OpenStudy (anonymous):

m is slope . You see that after the x :) It s 4/9

OpenStudy (anonymous):

These are the answer choices I don't see where u are coming from

OpenStudy (anonymous):

y = mx + c Gradient of the line acc to the law i stated = -2.25 y=-2.25x +c Substituting the co-ordinates 3=-2.25*4 +c 3=-9+c c=11 Therefore equation of the line is y=-2.25x + 12

OpenStudy (anonymous):

So the third option is correct acc to me What do u think @E.ali

OpenStudy (anonymous):

In Statistical Mathematics the product of the gradients of two lines is -1

OpenStudy (anonymous):